OpenStudy (keeli_grace):

Which is the quadratic variation equation for the relationship? y varies directly with x^2 and y = 72 when x = 6. A. y = 2x^2 B. y = 4x^2 C. y = 12x D. y = x^2 + 25

OpenStudy (keeli_grace):

@ParthKohli , could you please help me? :)

Parth (parthkohli):

\[y = kx^2\]The above is what a quadratic variation looks like. \(k\) is the constant of variation. Let's try to find out the constant of variation. It is given that when \(y =72\), \(x = 6\). Let's plug that in to our equation.\[y = kx^2 \implies 72 = k(6)^2\]

OpenStudy (keeli_grace):

I'm sorry, I'm absolutely terrible at math. Do you think you could give me a two-second math course? :/

Parth (parthkohli):

Well, I'm sorry about that. Do you know what a quadratic variation really is?

OpenStudy (keeli_grace):

Um, no.. :(

OpenStudy (anonymous):

@ParthKohli will help :)

OpenStudy (keeli_grace):

But I am a quick learner, so I'm sure I could figure it out fairly quickly if you helped me to understand! :)

Parth (parthkohli):

A "variation" is a relation between two quantities. It's like a pattern in how one quantity is, when you play around with another. If you look at a circle, the radius of the circle and the area are related. There is an equation between the two:\[\rm area = \pi \times (radius)^2\]If you make a new circle with twice the radius than one circle, then the new circle's area would not be twice the original circle's area: it would be FOUR times the first circle's area! If you make the radius thrice, the area will become NINE times the original. This is an example of a quadratic variation. When you make one quantity in a quadratic equation \(n\) times its original, the dependent quantity would be \(n^2\) times the original one. You can see why this happens from the equation of a variation.

OpenStudy (keeli_grace):

Okay, so far so good. You seem very passionate about algebra. :)

mathslover (mathslover):

Remember this : Parth is the great young Mathematician. Feel lucky to get help from him.

Parth (parthkohli):

Now, now, now - if \(y\) is a quantity that depends on \(x^2\), then as you increase \(x^2\), the more you will see an increase in \(y\). It is obvious why \(y = x^2\) is a quadratic variation. Now, for example, see \(y = 2x^2\). You will wonder why this is a quadratic variation at first. We saw how, from my last article, that if you multiply \(x\) by \(n\) in a variation, you will see that your new \(y\) is \(n^2\) times the original one. Let's try to see if the same happens in this case.\[y = 2x^2\]Multiply \(x\) by \(n\). The new quantity is \(nx\). Plug that for \(x\) in \(y = 2x^2\). You will get \(\text{the new }y = 2(nx)^2 \) by plugging in \(nx\) for \(x\). This simplifies to \[\text{the new y } = 2n^2 x^2 \\ ~~~~~~~ \quad ~~\quad = n^2 \times 2x^2 = n^2 \times y\]The new quantity \(y\) is \(n^2\) times the old \(y\) quantity.

Parth (parthkohli):

You can now see why this is called a "quadratic relationship" - quadratic literally means "the power two." Now, as I hope, you're pretty clear about what a quadratic relationship is. Let's now try to analyze many elements. Before I start, remind yourself that a quadratic variation is in the *form* \(y = kx^2\). It is not \(y = x^2\) itself. This means that all of the variations such as \(p^2 = 4q\) or \(\rm area = \pi r^2\) etc. are quadratic variations even though they don't initially appear to be. - - - ○ What is the difference between \(p = x^2\) and \(q = 2x^2\)? A: Let's try to see how this works. If you plug in \(x=1\) in both of the variations, we see that \(p = 1\) and \(q = 2\). When \(x = 2\), we see \(p = 4\) and \(q = 8\). If you continue, you will see how \(q\) is always twice \(p\). This means that \(q\) will be twice \(p\) for all \(x.\) We can extend this logic to variations \(y_1 = x^2\) and \(y_2 = kx^2\). For all \(x\), \(y_2 \) is \(k\) times \(y_1.\)

Parth (parthkohli):

The point I stated above is very, very obvious, but I felt a need to state it. --- Anyway. Now, when you have questions that tell you about a direct variation between two quantities \(y\) and \(x^2\), always try to see the information you have. Also remember that if you are given a direct variation, \(k\) represents a constant. If you know a value of \(k\), the same value of \(k \) will apply to all \(x\) and \(y\).

Parth (parthkohli):

One more thing: If \(x\) is related with \(y\), then \(y\) is related with \(k\). (Stating the obvious!) However, there is one small difference. When you are given \(y = k\cdot x\), then \(x = \frac{1}{k} y\). If \(x\) varies with \(y\) by a constant \(k\), then \(y\) varies with \(x\) by a constant \(\frac{1}k{}\).

Parth (parthkohli):

From my last post, just note one thing: if \(k\) is a constant, \(\frac{1}{k}\) is also a constant. Basically, to sum it all up, if \(\rm blah\) is in direct relation with \(\rm meh\), \(\rm bleh = k \cdot meh\). - - - - - Practice problem: \(y\) is directly related to \(2x^3\). When \(y = 32\), \(x\) is found to be \(1\). What is the constant of variation?